# m m OTr_ #1#3(#2#3)

### v1

@@ I Introduction
Thus, the conceptual gap between classical and quantum theories should be significantly reduced. However, quantum theory features incompatible observables that lead to effects unattainable by classical theory, even under an indeterministic interpretation thereof. We so identify the essence of quantum mechanics in phenomena involving Planck’s reduced constant, $\hbar$, that is, incompatible observables at the kinematical level. Heisenberg’s uncertainty relations, weak values [25], and the violation of Bell’s inequality, for example, rely on incompatible measurements interconnected by $\hbar$, and indeed lack classical analogues. It is of course very well-established that $\hbar$ characterizes quantum theory, but here we shed new light on the fact that, contrarily to the widespread belief, many features and paradoxes are not specific to quantum physics, unless they involve $\hbar$.
@@ IV A plurality of interpretations
Finally, interpretations like QBsim [46] (v) take the quantum state to be a subjective piece of information used by an agent to take rational decisions about what to “bet” on. Obviously, one can take such a subjective approach also in an indeterministic interpretation of classical physics. In fact, this simply amounts to acknowledge that there is no further underlying reality about which we assign probability. In this case one can still interpret physics in terms of FIQs, but what we have taken as propensities would be interpreted as the only available piece of information, i.e. epistemic probabilities. Therefore, what we have considered so far an objective process of actualization of propensities would be a Bayesian update of the information available to an agent.
@@ VI No-cloning
It was already shown that this result is not characteristic of the quantum formalism per se, but it holds unchanged for statistical ensembles of classical systems [53]. Let us recap the proof provided in Ref. [53].
@@ VIII Conclusions: what is really quantum is incompatibility
In essence, it appears to us that those phenomena rooted in the existence of incompatible variables, i.e., on the commutation relations involving the constant of nature $\hbar$, cannot be found in classical (indeterministic) theory. This is because in the latter we have only imposed that at any instant in time the state of a system has a finite volume $Vol(S)$, in phase space. However, in that case there is no relation whatsoever between the canonically conjugated variables appearing in phase space (e.g., position and momentum). In quantum theory, on the contrary, there is a smallest size for the volume (for each degree of freedom), i.e., $Vol(R)\geq(2\pi\hbar)^{\#(d.o.f.)}$, which establishes a relation between position and momentum. So, what really characterizes genuine quantum phenomena – at least at the kinematical level – is the Planck (reduced) constant $\hbar$. The dependence on $\hbar$ of genuine quantum features is manifest in uncertainty relations, [^10] In Ref. [1], it was shown that some form of uncertainty relations can be retrieved also in particular classical toy models (i.e., generalized-noncontextual ontological models). It is not clear how the latter relate to an alternative indeterministic classical physics – like the one based on FIQs– which do not establish any relation about the indeterminacy between conjugated variables. Some classical analogue of uncertainty relation is also discussed in Ref. [60] due to Gromov’s non-squeezing theorem in symplectic geometry. wave-particle duality [61], or weak values [25]. But this is the case also where it does not appear so explicitly, like in the violation of Bell inequalities which indeed require incompatible measurements [62], as well as in quantum key distribution [63], to name but a few examples.
@@ acknowledgements
We thank Johannes Fankhauser for useful discussions and for pointing out relevant literature. This research was supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699), and the Swiss National Science Foundation via the NCCR-SwissMap.

### v2

@@ I Introduction
Thus, the conceptual gap between classical and quantum theories should be significantly reduced. However, quantum theory features incompatible observables that lead to effects unattainable by classical theory, even under an indeterministic interpretation thereof. In general terms, incompatibility refers to pairs of physical variables that cannot be simultaneously determined with arbitrary precision. When one physical variable becomes more defined (for instance, due to a measurement), this generally affects the (in)determinacy of all variables incompatible with it. In quantum mechanics, incompatibility is associated with non-commuting observables, a relationship encapsulated by the Heisenberg uncertainty principle: the product of their indeterminacies (i.e., standard deviations) is bounded by a fundamental lower limit specific to the observables in question, in proportion to Planck’s reduced constant $\hbar$. For example, in the case of position and momentum, this lower bound is /2Planck-constant-over-$\hbar/2$. The situation is drastically different in classical theory, even in an indeterministic interpretation thereof, where states in phase space exhibit ontic indeterminacy. In fact, there is no posited lower limit to the joint determinacy that variables can acquire (again due to, e.g., measurements). In particular, position and momentum in phase space can be increasingly determined independently of each other, up to arbitrary, though always finite, precision.
We so identify the essence of quantum mechanics in phenomena involving Planck’s reduced constant, $\hbar$, that is, incompatible observables at the kinematical level. Heisenberg’s uncertainty relations, weak values [25], and the violation of Bell’s inequality, for example, rely on incompatible measurements interconnected by $\hbar$, and indeed lack classical analogues. It is of course very well-established that $\hbar$ characterizes quantum theory, but here we shed new light on the fact that, contrarily to the widespread belief, many features and paradoxes are not specific to quantum physics, unless they involve $\hbar$.
@@ IV A plurality of interpretations
Finally, interpretations like QBism [46] (v) take the quantum state to be a subjective piece of information used by an agent to take rational decisions about what to “bet” on. Obviously, one can take such a subjective approach also in an indeterministic interpretation of classical physics. In fact, this simply amounts to acknowledge that there is no further underlying reality about which we assign probability. In this case one can still interpret physics in terms of FIQs, but what we have taken as propensities would be interpreted as the only available piece of information, i.e. epistemic probabilities. Therefore, what we have considered so far an objective process of actualization of propensities would be a Bayesian update of the information available to an agent.
@@ VI No-cloning
It was already shown that this result is not characteristic of the quantum formalism per se, but it holds unchanged for statistical ensembles of classical systems [53, 54]. Let us recap the proof provided in Ref. [54].
@@ VIII Conclusions: what is really quantum is incompatibility
In essence, it appears to us that those phenomena rooted in the existence of incompatible variables, i.e., on the commutation relations involving the constant of nature $\hbar$, cannot be found in classical (indeterministic) theory. This is because in the latter we have only imposed that at any instant in time the state of a system has a finite volume $Vol(S)$, in phase space. However, in that case there is no relation whatsoever between the canonically conjugated variables appearing in phase space (e.g., position and momentum). In quantum theory, on the contrary, there is a smallest size for the volume (for each degree of freedom), i.e., $Vol(R)\geq(2\pi\hbar)^{\#(d.o.f.)}$, which establishes a relation between the (in)determinacy of position and momentum, a feature absent in classical theory. [^10] A partial classical analogue of uncertainty relations is discussed in Ref. [61], in the context of Gromov’s non-squeezing theorem in symplectic geometry. So, what really characterizes genuine quantum phenomena – at least at the kinematical level – is the Planck (reduced) constant $\hbar$. The dependence on $\hbar$ of genuine quantum features is manifest in uncertainty relations, [^11] In Ref. [1], it was shown that certain forms of uncertainty relations can also be retrieved in specific classical toy models (i.e., generalized non-contextual ontological models). However, we notice that the latter is an artificial toy model built to mimic certain features of quantum theory, such as incompatibility, while still adhering to some notion of classicality. The approach followed here, instead, is to take the actual theory of classical physics and relax its assumption of infinite precision, which leads to indeterminism without establishing any relation between the indeterminacy of conjugate variables. wave-particle duality [62], or weak values [25]. But this is the case also where it does not appear so explicitly, like in the violation of Bell inequalities which indeed require incompatible measurements [63], as well as in quantum key distribution [64], to name but a few examples.
@@ acknowledgements
We thank Johannes Fankhauser, Adan Cabello, Ämin Baumeler, and Mark Neyrinck for useful discussions and for pointing out relevant literature. This research was supported by the FWF (Austrian Science Fund) through an Erwin Schrödinger Fellowship (Project J 4699), and the Swiss National Science Foundation via the NCCR-SwissMap.